Leaf Industries is preparing its master budget for 2013. Relevant data pertaining to its sales budget are as follows:

Sales for the year are expected to total 8,000,000 units. Quarterly sales are 25%, 30%, 15%, and 30%, respectively. The sales price is expected to be $2.00 per unit for the first quarter and then be increased to $2.20 per unit in the second quarter.

1st Quarter

2nd Quarter

3rd Quarter

4th Quarter

Total for Yr

Unit Sales

8,000,000

Unit Selling Price

Total Sales
Answer #1

1st Quarter

2nd Quarter

3rd Quarter

4th Quarter

Total for Yr

Unit Sales

2,000,000 2,400,000 1,200,000 2,400,000

8,000,000

Unit Selling Price

$2.00 $2.20 $2.20 $2.20 $2.15

Total Sales

$4,000,000 $5,280,000 $2,640,000 $5,280,000

$17,200,000

1st Quarter Unit Sales = 0.25 *8,000,000= 2,000,000 units

2nd Quarter Unit Sales = 0.30 *8,000,000= 2,400,000 units

3rd Quarter Unit Sales = 0.15 *8,000,000= 1,200,000 units

4th Quarter Unit Sales = 0.30 *8,000,000= 2,400,000 units

Total Sales for each quarter can be calculated by multiplying Price in each Quarter with Unit Sales

Total Sales for Year=Sum of Total Sales for each Quarter = $17,200,000

Unit Selling Price for Year= $17,200,000/8,000,000 = $2.15
Answer #2
1st Quarter 2nd Quarter 3rd Quarter 4th Quarter Year
Unit Sales(a) 2,000,000 2,400,000 1,200,000 2,400,000 8,000,000
Unit Selling Price (b) 2.00 2.20 2.41 2.66 2.32
Total Sales 4,000,000 5,280,000 2,904,000 6,384,000 18,568,000

Note:

Unit Sales Calculation:

Q1- 8,000,000*0.25

Q2-8,000,000*0.30

Q3-8,000,000*0.15

Q4-8,000,000*0.30

Unit selling price is expected to grow at a rate of 10%, this is indicative from increased unit sale price per unit in Q2 from Q1.

Growth- 2.20/2.00 =1.1

Q3- 2.20*1.1 =2.42

Q4-2.42*1.1=2.66

Year- (2.00+2.20+2.42+2.66)/4 =2.32
Answer #3

Let x be quarterly sales

So to find value of x :-

25x + 30x + 15x + 30x = 8,000,000

100x = 8,000,000

X = 80,000 units

1stQuarter

2ndQuarter

3rdQuarter

4thQuarter

Total for Yr

Unit Sales

2,000,000

( 80000 * 25)

2,400,000

(80000 *30)

1,200,000

(80000 * 15)

2,400,000

(80000* 30)

8,000,000

Unit Selling Price

$ 2.00

$ 2.20

$ 2.20

$ 2.20

Total Sales

4,000,000

5,280,000

2,640,000

5,280,000

6,640,000